# A Bayesian Estimator of Sample Size

### v1

@@ 4.1 Correlation between sample size, evidence, and confidence
Given the discrete nature of the outcomes, it’s not always possible to find an evidence level for each sample size $n$ that exactly matches $\text{Pr}(H=H_{1}|e,n)=0.6$. Thus, we document the smallest $e$ for each $n$ where $\text{Pr}(H=H_{1}|e,n)\geq 0.6$. In this case, for $\theta^{*}=0.05$, we report $n(e)$, the different sample sizes $n$ with their corresponding evidence $e$: $50(0.080)$, $100(0.070)$, $150(0.067)$, $200(0.065)$, and $1000(0.057)$. These values clearly show that as the sample size increases, the evidence needed to maintain the confidence at 0.6 decreases, approaching $\theta^{*}=0.05$. Similar results can be obtained for continuous and count-data outcomes in two-arm trial, also the three outcomes in one-arm trial.
@@ 4.2 Coherence between BESS and Bayesian Inference
After the trial is conducted and data is observed, a Bayesian analysis of the observed data using posterior probability will be coherent with the BESS statement. To explain this, we first denote $\bm{y}^{\star}_{n}$ the observed data after the trial is completed, $e^{\star}$ the observed evidence, and $c^{\star}=\text{Pr}(H_{1}|\bm{y}^{\star}_{n})$ the posterior probability of $H_{1}$ conditional on the observed data $\bm{y}^{\star}_{n}.$ Then the coherence between BESS and Bayesian inference means that if $e^{\star}>e$, $c^{\star}>c.$
This type of coherence is important for BESS to be adopted in practice since investigators of clinical trials are usually not statisticians, and the coherence property of BESS allows them to connect the design (i.e., sample size statement) of the trial with the statistical analysis of the observed data once the trial is carried out.
To verify the coherence numerically, we perform a simple simulation. In Appendix Table A.1 we set up the true parameters for data simulation for each type of clinical trials. For example, for two-arm binary data, we assume the minimum treatment effect $\theta^{*}=0.05$, evidence $e=0.1$, confidence $c=0.8$. Using a prior distribution $\theta_{j}\sim\text{Beta}(0.5,0.5)$ BESS leads to a sample size of 150. We then repeatedly generate trial data assuming different true values of $\theta_{1}$ and $\theta_{0}$ in Table A.1, and report the inference results of $e^{\star}$ and $c^{\star}$. Figure 2 summarizes the simulation results. In all three subplots (a)-(c), when $e^{\star}$ is larger than $e$ (the black vertical line), $c^{\star}$ is larger than $c$, the red horizontal line.
@@ 5.1 Simulation Setup
Through simulation, we compare BESS and standard SSE for a two-arm trial with binary outcomes. For fair comparison, we match the Type I/II error rates for both methods. The matching is realized by three steps. Step 1: we obtain sample size estimated via the proposed BESS approach for a trial. Step 2: we repeatedly simulate trial data under the null and alternative hypotheses, perform the Bayesian inference using the same model as in BESS, and record the Type I and Type II error rates from the simulated trials. Step 3: using the two error rates, we apply the standard SSE to arrive at a Frequentist sample size, and compare it with the BESS estimate. See Appendix Figure A.1 for an illustration.
@@
Based on a set of $\theta_{1}$ and $\theta_{0}$ values, we generate the binary responses of 150 patients each in the treatment arm and control arm. Denote the simulated outcomes $\bm{y}_{n}=\{\bm{y}_{n1}\text{ and }\bm{y}_{n0}\}$, $\bm{y}_{nj}=\{y_{ij};i=1,\ldots,n\},\,j=0,1.$ We generate $y_{ij}\sim\text{Bern}(\theta_{j})$ where $\text{Bern}(\theta)$ is a Bernoulli distribution with mean $\theta.$ The null is rejected if $\text{Pr}(H=H_{1}|\bm{y}_{n})>0.8.$ Here, $0.8$ is arbitrarily selected without specific intention. A larger or smaller cutoff than $0.8$ arbitrarily affects the computed Type I/II error rates, which does not affect the objective of our simulation. We simulate 10,000 trials, each under the null and alternative. A rejection of the null for a trial simulated under the null is recorded as an incidence of Type I error and a non-rejection of the null for a trial under the alternative is recorded as an incidence of Type II error. The Type I/II error rates $(\alpha/\beta)$ are then computed as the frequencies of the corresponding incidences over the 10,000 trials.
@@
Based on the computed Type I/II error rates $(\alpha/\beta)$ from Step 2, denoted as $\alpha$ and $\beta$, we estimate a sample size using a standard SSE approach. For example, we consider a superiority $z$-test for comparing $\theta_{1}$ and $\theta_{0}$, and a sample size can be estiamted via (6). In the estimation, we assume the true values for $\theta_{1}$ and $\theta_{0}$, which gives the standard approach an “oracle” performance. In other words, the estimated sample size is guaranteed to achieve the target Type I/II error rates $\alpha$ and $\beta$, since the assumed $\theta_{1}$ and $\theta_{0}$ in the sample size estimation match the true values. Even though this is typically not achievable in reality, we decide to compare the oracle Frequentist sample size with the BESS sample size nevertheless.
@@ 5.2 Simulation Result
Several trends are worth noting in Table 2. First, increase confidence cutoff $c$ leads to lower Type I error rate for fixed evidence $e.$ This is because increase in cutoff $c$ makes it harder to reject the null for BESS under (7), and hence fewer simulated trials under $H_{0}$ will be rejected, i.e., decrease in the Type I error rate. Second, note that the true effect size under $H_{1}$ equals $\theta\equiv\theta_{1}-\theta_{0}=0.4-0.25=0.15.$ We observe that 1) power increases when the confidence cutoff $c$ increases and evidence $e$ is less than the the true effect size, i.e., $e=0.1,$ 2) power stays the same if evidence $e$ equals the true effect size, and 3) power decreases when confidence increases and $e$ is greater than the true effect size. This complicated trend demonstrates an interaction between $c$ and power conditional on whether $e$ is greater than, equal to, or less than the true effect size $\theta$. To understand this, first recall in Section 4.2 we used notation $e^{\star}$ to denote the observed evidence after the trial is completed and data observed. When $c$ increases, the BESS-estimated sample size also increases (see Section 4.1). Therefore, the observed $e^{\star}$ will be closer to $\theta$ due to large number theory. Consequently, more simulated trials under the alternative will see $e^{\star}$ close to $\theta$. If the assumed evidence $e$ is smaller than $\theta$, more likely it will be smaller than $e^{\star}$ as well. This means that the posterior probability $\text{Pr}(H_{1}|\bm{y}_{n})$ will be higher (since there is stronger evidence supporting $H_{1}$), and hence more rejections, i.e., higher power. Therefore, when $e$ is smaller than $\theta$, a larger $c$ leads to higher power. Same logic applies to the case when $e$ is larger than $\theta,$ in which a larger $c$ leads to lower power. Lastly, when $e=\theta$, increasing sample size (as a result of increasing $c$) makes $e^{\star}$ approach $e$, and therefore there is no obvious impact on the power.
@@
Finally, from a true Bayesian perspective, BESS is concerned about the trial at hand, rather than hypothetical trials generated from the null or alternative. Therefore, while Table 2 illustrates a connection between BESS and standard SSE, it does not imply that BESS needs to be calibrated based on Type I/II error rates in practice. On the contrary, BESS focuses on the probability of making a right decision given the observed data, which can be measured by the false positive rate (FPR) and false negative rate (FNR). In Table 2 the reported FPR and FNR for BESS and standard SSE are the same since 1) the prevalence of trials under the $H_{0}$ and $H_{1}$ is 50% and 2) the standard SSE is oracle since it assumes the true $\theta_{0}$ and $\theta_{1}$ values in its computation. In Appendix Table A.2 we show that when the $\theta_{0}$ and $\theta_{1}$ are mis-specified in the standard SSE, the estimated sample sizes may be too large or too small, leading to over- or under-power, and deflated or inflated FPR/FNR’s.
@@
We demonstrate the sensitivity of incorporating prior information through simulation, assuming there exists prior data of $n_{0}$ patients per arm. The simulation details are presented in Appendix 5.1 , and the average sample size from the 1,000 simulated trials is 31.42 with a standard deviation of 28. Recall with a Beta(0,0)00(0,0)( 0 , 0 ) prior the BESS sample size was 40 in Table 2 for $e=0.15$ and $c=0.8$. The results show that BESS is able to properly borrow prior information for sample size estimation. In practice, one may use different informative priors, e.g, power prior (Ibrahim and Chen, 2000) or commensurate prior (Hobbs et al., 2011), for information borrowing. We leave these topics for future research.
@@ 6.1 Fixed Sample Size
BESS provides the following sample size statement: Assuming evidence $e=0$, a sample size of 57 patients per arm is needed to declare with 70% confidence that the response rate of the higher dose is no higher than the lower dose by 0.05.
@@ 1. BESS SSR
In the SSR, we again use BESS to re-estimate the sample size based on updated evidence from the interim data, defined as $e_{\text{int}}=\text{E}[\theta_{L}|\bm{y}_{n/2}]-\text{E}[\theta_{H}|\bm{y}_{n/2}]$, which corresponds to the difference of posterior means. We use the same cutoff $c$ for the SSR. However, we use the posterior distribution $\pi(\theta_{L},\theta_{H}|\bm{y}_{n/2})$ as the prior in the SSR using the BESS algorithm 2. Denote the additional sample size as $n^{*}$ based on SSR. At the end of the trial after $n^{*}$ more patients are randomized to each dose, BESS SSR rejects the null and accepts the alternative if $\text{Pr}(H_{1}|\bm{y}_{n/2+n^{*}})\geq c$. Otherwise, it accepts the null and rejects the alternative.
@@ 4. Standard SSE with interim
We assume $q=0.5$. For Designs 1 & 2, we let $e=0.02,$ $c=0.7$ and $c^{*}=0.3$. For designs 3 & 4, we set $\alpha=0.3$ and $\beta=0.3$, so that BESS and standard SSE produce the same sample size for the trial in the beginning, which is $n=100.$ Lastly, we assume $(\theta_{L},\theta_{H})\sim\text{Beta}(0.05,0.05)\text{Beta}(0.05,0.05)I(\theta\in H)$, where $\text{Beta}(0.05,0.05)$ is chosen to reduce the prior effective sample size. A total of 2,000 simulated trials are generated, 1,000 each under $H_{0}$ or $H_{1}$, for each design in each scenario. Four metrics are used to compare the designs, Type I/II error rates, false positive rate (FPR) and false negative rate (FNR). See Appendix A.7 for detail.
@@ Results
where the weight $k\in[0,\infty)$ is a pre-determined factor that quantifies the relative weight between the Type I and Type II error rates. We let $k\in\{0.5,1,1.5\}$. A lower CER is more desirable. Top two rows in Figure 3 reports the results for the four designs. Across all designs, as the sample size increases, the rate at which CER declines, indicating potential diminishing return when sample size gets larger. In other words, the gain in reduction of error rates may lessen when sample size continues to increase. Considering the substantial costs associated with patient enrollment, these findings suggest finding a “sweet spot” of sample size that achieves a desirable tradeoff between cost and statistical properties. Comparing across designs, we find that except for the Standard SSE design, the performance of the other designs are similar. Notably, the BESS SSR Cap design (design 2) seems to be in general the best across most cases. In Scenario 2 when $k=0.5$, the standard SSE is the winning design. In contrast, it is the losing design in the same scenario but with $k=1.5.$ This seems to suggest the standard SSE is a better design if the Frequentist Type I error rate is of important consideration in design evaluation. However, in early-phase dose optimization trials, Type I error rate is not the primary concern. In fact, one may argue it is of the least concern. For example, one would be much more concerned if a “GO” decision that recommend a wrong dose to further clinical development, or a “No Go” decision that fails to recommend a promising dose. These are measured by the FPR and FNR in the bottom two rows of Figure 3.
@@ 7 Discussion
Many future directions may be considered to further develop BESS and corresponding adaptive designs, such as sample size re-estimations, downgrade historical data in prior construction, decision rules for futility stopping, etc.
@@ A.4 BESS Algorithm 2’
A.6 Additional Simulation Setup and Results in Section 5.1

### v2

@@ 4.1 Correlation between sample size, evidence, and confidence
Given the discrete nature of the outcomes, it’s not always possible to find an evidence level for each sample size $n$ that exactly matches $\text{Pr}(H=H_{1}|e,n)=0.6$. Thus, we document the smallest $e$ for each $n$ where $\text{Pr}(H=H_{1}|e,n)\geq 0.6$. In this case, for $\theta^{*}=0.05$, we report $n(e)$, the different sample sizes $n$ with their corresponding evidence $e$: $50(0.080)$, $100(0.070)$, $150(0.067)$, $200(0.065)$, and $1000(0.057)$. These values clearly show that as sample size increases, evidence needed to maintain the confidence at 0.6 decreases, approaching $\theta^{*}=0.05$.
@@ 4.2 Coherence between BESS and Bayesian Inference
After the trial is conducted and data observed, a Bayesian analysis of the observed data using posterior probability will be coherent with the BESS statement. To explain this, we first denote $\bm{y}^{\star}_{n}$ the observed data after the trial is completed, $e^{\star}$ the observed evidence, and $c^{\star}=\text{Pr}(H_{1}|\bm{y}^{\star}_{n})$ the posterior probability of $H_{1}$ conditional on the observed data $\bm{y}^{\star}_{n}.$ Then the coherence between BESS and Bayesian inference means:
If $e^{\star}>e$, then $c^{\star}>c.$
Coherence is important for BESS to be adopted in practice since investigators of clinical trials are usually not statisticians, and the coherence property of BESS allows them to connect the design (i.e., sample size statement) of the trial with the statistical analysis of the observed data once the trial is carried out.
To verify the coherence property numerically, we perform a simple simulation. In Appendix Table A.1 we set up the true parameters for data simulation for each type of clinical trial data. For example, for two-arm binary data, we assume the minimum treatment effect $\theta^{*}=0.05$, evidence $e=0.1$, confidence $c=0.8$. Using a prior distribution $\theta_{j}\sim\text{Beta}(0.5,0.5)$ BESS leads to a sample size of 150. We then repeatedly generate trial data assuming different true values of $\theta_{1}$ and $\theta_{0}$ in Table A.1, and report the inference results of $e^{\star}$ and $c^{\star}$. Figure 2 summarizes the simulation results. In all three subplots (a)-(c), when $e^{\star}$ is larger than $e$ represented by the black vertical line, $c^{\star}$ is larger than $c$, the red horizontal line.
@@ 5.1 Simulation Setup
Through simulation, we compare BESS and the standard SSE for a two-arm trial with binary outcome. For fair comparison, we match the type I/II error rates for both methods. The matching is realized by three steps. Step 1: we obtain sample size estimated via the proposed BESS approach for a trial. Step 2: we repeatedly simulate trial data under the null and alternative hypotheses, perform the Bayesian inference using the same model as in BESS, and record the type I and type II error rates from the simulated trials. Step 3: using the two error rates, we apply the standard SSE to arrive at a Frequentist sample size, and compare it with the BESS estimate. See Appendix Figure A.1 for an illustration.
@@
Based on each pair of $\theta_{1}$ and $\theta_{0}$ values, we generate the binary responses of 150 patients per arm for the treatment arm and control arm. Denote the simulated outcomes $\bm{y}_{n}=\{\bm{y}_{n1}\text{ and }\bm{y}_{n0}\}$, $\bm{y}_{nj}=\{y_{ij};i=1,\ldots,n\},\,j=0,1.$ We generate $y_{ij}\sim\text{Bern}(\theta_{j})$ where $\text{Bern}(\theta)$ is a Bernoulli distribution with mean $\theta.$ The null is rejected if $\text{Pr}(H=H_{1}|\bm{y}_{n})>0.8.$ Here, $0.8$ is arbitrarily selected without specific intention. A larger or smaller cutoff than $0.8$ arbitrarily affects the computed type I/II error rates, which does not affect the objective of our simulation. We simulate 10,000 trials each under the null and alternative. A rejection of the null for a trial simulated under the null is recorded as an incidence of type I error and a non-rejection of the null for a trial under the alternative is recorded as an incidence of type II error. The type I/II error rates $(\alpha/\beta)$ are then computed as the frequencies of the corresponding incidences over the 10,000 trials.
@@
Based on the computed type I/II error rates $(\alpha/\beta)$ from Step 2, denoted as $\alpha$ and $\beta$, we estimate a sample size using a standard SSE approach. For example, we consider a superiority $z$-test for comparing $\theta_{1}$ and $\theta_{0}$, and a sample size can be estiamted via (6). In the estimation, we assume the true values for $\theta_{1}$ and $\theta_{0}$ under the alternative, which gives the standard approach an “oracle” performance. In other words, the estimated sample size is guaranteed to achieve the target Type I/II error rates $\alpha$ and $\beta$, since the assumed $\theta_{1}$ and $\theta_{0}$ in the sample size estimation match the true values. Even though this is typically not achievable in reality, we decide to compare the oracle Frequentist sample size with the BESS sample size nevertheless.
@@ 5.2 Simulation Result
Several trends are worth noting in Table 2. First, increasing confidence cutoff $c$ leads to lower type I error rate for fixed evidence $e.$ This is because increase in cutoff $c$ makes it harder to reject the null for BESS under (7), and hence fewer simulated trials under $H_{0}$ will be rejected, i.e., decrease in the type I error rate. Second, note that the true effect size under $H_{1}$ equals $\theta\equiv\theta_{1}-\theta_{0}=0.4-0.25=0.15.$ We observe that 1) power increases when the confidence cutoff $c$ increases and evidence $e$ is less than the true effect size, i.e., $e=0.1,$ 2) power stays the same if evidence $e$ equals the true effect size, and 3) power decreases when confidence increases and $e$ is greater than the true effect size. This complicated trend demonstrates an interaction between $c$ and power conditional on whether $e$ is greater than, equal to, or less than the true effect size $\theta$. To understand this, first recall in Section 4.2 we used notation $e^{\star}$ to denote the observed evidence after the trial is completed and data observed. When $c$ increases, the BESS-estimated sample size also increases (see Section 4.1). Therefore, the observed $e^{\star}$ will be closer to $\theta$ due to large number theory. Consequently, more simulated trials under the alternative will see $e^{\star}$ close to $\theta$. If the assumed evidence $e$ is smaller than $\theta$, more likely it will be smaller than $e^{\star}$ as well. This means that the posterior probability $\text{Pr}(H_{1}|\bm{y}_{n})$ will be higher (since there is stronger evidence supporting $H_{1}$), and hence more rejections, i.e., higher power. Therefore, when $e$ is smaller than $\theta$, a larger $c$ leads to higher power. Same logic applies to the case when $e$ is larger than $\theta,$ in which a larger $c$ leads to lower power. Lastly, when $e=\theta$, increasing sample size (as a result of increasing $c$) makes $e^{\star}$ approach $e$, and therefore there is no obvious impact on the power.
@@
Finally, from a true Bayesian perspective, BESS is concerned about the trial at hand, rather than hypothetical trials generated from the null or alternative (i.e., type I/II error rates). Therefore, while Table 2 illustrates a connection between BESS and standard SSE, it does not imply that BESS needs to be calibrated based on type I/II error rates in practice. On the contrary, BESS focuses on the probability of making a right decision given the observed data, which can be measured by the FPR and FNR. In Table 2 the reported FPR and FNR for BESS and standard SSE are the same since 1) the prevalence of trials under the $H_{0}$ and $H_{1}$ is 50% and 2) the standard SSE is oracle since it assumes the true $\theta_{0}$ and $\theta_{1}$ values in its computation. In Appendix Table A.2 we show that when the $\theta_{0}$ and $\theta_{1}$ are mis-specified in the standard SSE, the estimated sample sizes may be too large or too small, leading to over- or under-power, and deflated or inflated FPR/FNR’s.
@@
We demonstrate the sensitivity of incorporating prior information through simulation, assuming there exists prior data of $n_{0}$ patients per arm. The simulation details are presented in Appendix A.6, and the average sample size from the 1,000 simulated trials is 31.42 with a standard deviation of 28. Recall with a Beta(0,0)00(0,0)( 0 , 0 ) prior the BESS sample size was 40 in Table 2 for $e=0.15$ and $c=0.8$. The results show that BESS is able to properly borrow prior information for sample size estimation. In practice, one may use different informative priors, e.g, power prior (Ibrahim and Chen, 2000) or commensurate prior (Hobbs et al., 2011), for information borrowing. We leave these topics for future research.
@@ 6.1 Fixed Sample Size
BESS provides the following sample size statement: Assuming evidence $e=0$, a sample size of 57 patients per arm is needed to declare with 70% confidence that the response rate of the higher dose is no higher than the lower dose by 0.05. Larger or smaller sample size can be achieved by calibrating $e$ and $c$.
@@ 1. BESS SSR
In the SSR, we again use BESS to re-estimate the sample size based on updated evidence from the interim data, defined as $e_{\text{int}}=\text{E}[\theta_{L}|\bm{y}_{n/2}]-\text{E}[\theta_{H}|\bm{y}_{n/2}]$, which corresponds to the difference of posterior means. We use the same cutoff $c$ for the SSR. However, we use the posterior distribution $\pi(\theta_{L},\theta_{H}|\bm{y}_{n/2})$ as the prior in the SSR applying the BESS algorithm 2. Denote the additional sample size as $n^{*}$ based on SSR. At the end of the trial after $n^{*}$ more patients are randomized to each dose, BESS SSR rejects the null and accepts the alternative if $\text{Pr}(H_{1}|\bm{y}_{n/2+n^{*}})\geq c$. Otherwise, it accepts the null and rejects the alternative.
@@ 4. Standard SSE with interim
We assume $q=0.5$. For Designs 1 & 2, we let $e=0.02,$ $c=0.7$ and $c^{*}=0.3$. For designs 3 & 4, we set $\alpha=0.3$ and $\beta=0.3$, so that BESS and standard SSE produce the same sample size for the trial in the beginning, which is $n=100.$ Lastly, we assume $(\theta_{L},\theta_{H})\sim\text{Beta}(0.05,0.05)\text{Beta}(0.05,0.05)I(\theta\in H)$, where $\text{Beta}(0.05,0.05)$ is chosen to reduce the prior effective sample size. A total of 2,000 simulated trials are generated, 1,000 each under $H_{0}$ or $H_{1}$, for each design in each scenario.
@@ Results
where the weight $k\in[0,\infty)$ is a pre-determined factor that quantifies the relative weight between the type I and type II error rates. We let $k\in\{0.5,1,1.5\}$. A lower CER is more desirable. Top two rows in Figure 3 reports the results for the four designs. Across all designs, as the sample size increases, the rate at which CER declines, indicating potential diminishing return when sample size gets larger. In other words, the gain in reduction of error rates may lessen when sample size continues to increase. Considering the substantial costs associated with patient enrollment, these findings suggest finding a “sweet spot” of sample size that achieves a desirable tradeoff between cost and statistical properties. Comparing across designs, we find that except for the Standard SSE design, the performance of the other designs are similar for the same average sample size. Notably, the BESS SSR Cap design (design 2) seems to be in general the best across most cases as it shows slightly lower CER for fixed sample size. In Scenario 2 when $k=0.5$, the standard SSE is the winning design. In contrast, it is the losing design in the same scenario but with $k=1.5.$ This seems to suggest the standard SSE is a better design if the Frequentist type I error rate is of important consideration in design evaluation. However, in early-phase dose optimization trials, type I error rate is not the primary concern. In fact, one may argue it is of the least concern. For example, one would be much more concerned if a “GO” decision that recommend a wrong dose to further clinical development, or a “No Go” decision that fails to recommend a promising dose. These are measured by the FPR and FNR in the bottom two rows of Figure 3.
@@ 7 Discussion
Many future directions may be considered to further develop BESS and corresponding adaptive designs, such as group sequential designs with sample size re-estimations, and down-weight historical data in prior construction and decision rules for futility stopping, etc.
@@ A.4 BESS Algorithm 2’
A.6 Additional Simulation Setup and Results in Section 5